# The Picard group of the Baily-Borel compactification of the moduli space of polarized K3 surfaces

### v1

@@ Appendix A A vector-valued Siegel-Weil formula
we have $\omega_{\psi,f}(g_{f})\circ\iota=\iota\circ\overline{\rho_{M}^{(d)}}(g)$ (cf. [36]) for $g\in{\rm Mp}_{2d}(\mathbb{Z})$ and $g_{f}\in{\rm Mp}_{2d}(\hat{\mathbb{Z}})$ the unique element such that $gg_{f}\in{\mathrm{Sp}}_{2d}(\mathbb{Q})\subset{\rm Mp}_{2d}(\mathbb{A})$.
@@
$\theta(g,h,\varphi):=\sum_{\mathbf{x}\in V(\mathbb{Q})^{d}}\omega_{\psi}(g)\varphi(h^{-1}\mathbf{x}),\quad g\in{\rm Mp}_{2d}(\mathbb{A}),\quad h\in\mathrm{O}(V)(\mathbb{A}).$
@@
$A(R):=\left\{m(U):=\left(\begin{array}[]{cc}U&0\\ 0&(U^{-1})^{\top}\end{array}\right)|U\in{\rm GL}_{d}(R)\right\},$
@@
$N(R):=\left\{n(B):=\left(\begin{array}[]{cc}I_{d}&B\\ 0&I_{d}\end{array}\right)|B\in{\rm Sym}_{d}(R)\right\}.$
@@
$E(g,s,\varphi):=\sum\limits_{\gamma\in P(\mathbb{Q})\backslash{\mathrm{Sp}}_{2d}(\mathbb{Q})}\Phi(\gamma g,s,\varphi)\quad g\in{\rm Mp}_{2d}(\mathbb{A}),\quad s\in\mathbb{C},$
@@
The famous Siegel-Weil formula identifies the Theta lifting of the constant function $\mathbf{1}$ and the special value of the Eisenstein series.
Theorem A.4 (Siegel-Weil formula in the metaplectic case (cf. [34])).
@@
Suppose $M$ is a negative definite even lattice of rank $r$ with $\frac{r}{2}>d+1$. Then
@@
The idea is to identify the component functions of both sides via the Siegel-Weil formula. When $r$ is even, this is [25, Theorem 5.5]. Now suppose $r$ is odd and we sketch the proof as below.
According to Kudla (cf. [20, p. 37 Proposition 4.3]), for the archimedean component of the Weil representation $\omega_{\psi}$, there is a character $\chi_{V,\infty}^{\psi}$ of $\widetilde{A(\mathbb{R})}$ such that
$(\omega_{\psi,\infty}(\tilde{m})\varphi_{\infty})(\mathbf{x})=\chi_{V,\infty}^{\psi}(\tilde{m})|\det(m(a))|^{\frac{r}{2}}\varphi_{\infty}(\mathbf{x}\cdot a),\quad\tilde{m}=(m(a),\phi_{m(a)})\in\widetilde{A(\mathbb{R})},$
where $\varphi_{\infty}\in\mathcal{S}(V(\mathbb{R})^{d})$ and $a\in{\rm GL}_{d}(\mathbb{R})$.
Let $\varphi_{\infty}(\mathbf{x})=e^{-\pi\mathrm{tr}(\mathbf{x},\mathbf{x})}$ be the standard Gaussian function and for $\gamma\in(M^{\vee}/M)^{d}$, denote $\varphi_{\infty}\otimes\varphi_{\gamma}$ by $\varphi_{\infty,\gamma}$. For $\tau=x+iy\in\mathbb{H}_{d}$ and $a\in{\rm GL}_{d}(\mathbb{R})$ satisfying $aa^{\top}=y$, we consider the element
@@
The component function of our previously defined vector-valued genus Theta series (resp. vector-valued Siegel-Eisenstein series) is given as follows.
@@
    $\langle\Theta^{(d)}_{\mathbf{Gen}(M)}(\tau),\mathfrak{e}_{\gamma}\rangle=\chi_{V,\infty}^{\psi}(\tilde{m})^{-1}|\det(m(a))|^{-\frac{r}{2}}\int_{\mathrm{O}(V)(\mathbb{Q})\backslash\mathrm{O}(V)(\mathbb{A})}\theta(\widetilde{g}_{\tau},h,\varphi_{\infty,\gamma})\mathrm{d}h,$
    where $\widetilde{g}_{\tau}=n(x)\tilde{m}=(g_{\tau},\phi_{g_{\tau}})$ and the Haar measure $\mathrm{d}h$ is normalized so that ${\rm vol}(\mathrm{O}(V)(\mathbb{Q})\backslash\mathrm{O}(V)(\mathbb{A}))=1$.
  • Siegel-Eisenstein series:
@@
For the comparison of Theta series, the proof is exactly the same as in the classical case (cf. [23, Example 2.2.6]). For more details, one can see [25, Theorem 5.5]. For the comparison of Siegel-Eisenstein series. Since [25, Theorem 5.5] uses a computation by Kudla which might be difficult to locate a reference in the metaplectic case, we briefly present an alternative proof here.
Consider $q_{\tau}(\mathbf{x}):=e^{\pi i\mathrm{tr}((\mathbf{x},\mathbf{x})\tau)}$. One has
@@
for $n(B)=\Big{(}\left(\begin{array}[]{cc}\rm{I}_{d}&B\\ 0&\rm{I}_{d}\end{array}\right),1\Big{)}$, where $B\in{\rm Sym}_{d}(\mathbb{Z})$ and
@@
for $J_{d}=\Big{(}\left(\begin{array}[]{cc}0&-\rm{I}_{d}\\ \rm{I}_{d}&0\end{array}\right),\sqrt{\det(\tau)}\Big{)}$. Since ${\rm Mp}_{2d}(\mathbb{Z})$ is generated by $J_{d}$ and $n(B)$’s, combining these equations one gets
@@
for $\tilde{g}=(g,\phi_{g}(\tau))\in{\rm Mp}_{2d}(\mathbb{Z})$. Hence
@@
and the rest are exactly the same as the last several lines in [25, Theorem 5.5].
@@ Appendix B Coset decomposition and Weil representations
It has been shown in [25, Proposition 4.2] that for any $g^{(2)}\in{\mathrm{Sp}}_{4}(\mathbb{Z})$ there exist some $\alpha\in\mathbb{Z}^{\leq 0},A,B\in{\rm SL}_{2}(\mathbb{Z})$ such that
@@
and
$\varphi(D\mathcal{C}_{\alpha}u(A)d(B))=\pm B^{\prime}\begin{pmatrix}\alpha^{2}&0\\ 0&1\end{pmatrix}A,$
@@ 2. $(ii)$
$\phi_{(A\cdot B)^{\prime}}(z)\sqrt{z^{\prime}+(A\cdot B)^{\prime}z}=\phi_{A^{\prime}}(z)\phi_{B}(z^{\prime})\sqrt{A^{\prime}z+Bz^{\prime}}=\phi_{B^{\prime}\cdot A^{\prime}}(z)\sqrt{z^{\prime}+(A\cdot B)^{\prime}z}.$
@@
This is clear because
@@
Now it follows from [25, Proposition 4.2] that there is a well-defined injection
@@
Lemma B.3.
[25, Lemma 4.3] Let $\alpha\in\mathbb{Z}$, $\tilde{A}\in\widetilde{\mathbf{Y}}_{\alpha^{2}}$ and $\tilde{B}\in{\rm Mp}_{2}(\mathbb{Z})$. Then
@@
Given this decomposition, when $\tilde{g}^{(2)}=\widetilde{\mathcal{C}}_{\alpha}$, the equation B.5 holds by a direct computation (cf. [25, Proposition 4.4]). [^2] A careful reader may note that since our definition of ”$J$” is different from that in [25], the action of ”$J_{2}$” in [25] is actually the action of ”$J_{2}^{-1}$” in our notation. Therefore, the reason why we use $J_{2}^{-1}$ in the above decomposition is to keep the consistency with the computations in [25, Proposition 4.4]. Then from the decomposition (B.1), we have
$\rho_{M}^{(2)}(\tilde{g}^{(2)})^{-1}(\mathfrak{e}_{0}\otimes\mathfrak{e}_{0})$ $=\rho_{M}^{(2)}(\widetilde{\mathcal{C}}_{\alpha}u(\tilde{A})d(\tilde{B}))^{-1}(\mathfrak{e}_{0}\otimes\mathfrak{e}_{0})$ (B.6)
@@ Appendix C Hecke kernels
Denote the element $\big{(}T=\begin{pmatrix}1&1\\ 0&1\end{pmatrix},1\big{)}\in{\rm Mp}_{2}(\mathbb{Z})$ by $\widetilde{T}$. We have
$K_{1}(z,z^{\prime})$ $=\sum\limits_{\gamma\in G_{M}}\sum\limits_{\tilde{g}\in{\rm Mp}_{2}(\mathbb{Z})}\frac{\phi_{g}(z)^{-2k}}{(z^{\prime}+g\cdot z)^{k}}\rho_{M}(\tilde{g})^{-1}(\mathfrak{e}_{\gamma})\otimes\mathfrak{e}_{\gamma}$ (C.2)
@@
where $\mathbf{P}_{\gamma,m}(z)=\frac{1}{2}\sum\limits_{\tilde{g}\in\widetilde{\Gamma}_{\infty}^{+}\backslash{\rm Mp}_{2}(\mathbb{Z})}(\mathbf{e}(mz)\mathfrak{e}_{\gamma})\mid_{k,G_{M}}[\tilde{g}]$ is the Poincaré series defined in [10, §1.2]. Here the last identity comes from the Lipschitz Summation Formula in [24] (see also [28, Theorem 1]), which we recall in Lemma C.2 below.
@@
By [10, Proposition 1.5], one has $\langle f,\mathbf{P}_{\gamma,m}(z)\rangle_{\rm Pet}=\frac{2\cdot\Gamma(k-1)}{(4\pi m)^{k-1}}c(m,\gamma)$. Combine this with (C.2), we finally get

### v2

@@ Appendix A A vector-valued Siegel–Weil formula
we have $\omega_{\psi,f}(g_{f})\circ\iota=\iota\circ\overline{\rho_{M}^{(d)}}(g)$ (cf. [Zhang2009ModularityOG]) for $g\in{\rm Mp}_{2d}(\mathbb{Z})$ and $g_{f}\in{\rm Mp}_{2d}(\hat{\mathbb{Z}})$ the unique element such that $gg_{f}\in{\mathrm{Sp}}_{2d}(\mathbb{Q})\subset{\rm Mp}_{2d}(\mathbb{A})$.
@@
$\theta(g,h,\varphi)\coloneqq\sum_{\mathbf{x}\in V(\mathbb{Q})^{d}}\omega_{\psi}(g)\varphi(h^{-1}\mathbf{x}),\quad g\in{\rm Mp}_{2d}(\mathbb{A}),\quad h\in\mathrm{O}(V)(\mathbb{A}).$
@@
$A(R)\coloneqq\left\{m(U)\coloneqq\left(\begin{array}[]{cc}U&0\\ 0&(U^{-1})^{\top}\end{array}\right)|U\in{\rm GL}_{d}(R)\right\},$
@@
$N(R)\coloneqq\left\{n(B)\coloneqq\left(\begin{array}[]{cc}I_{d}&B\\ 0&I_{d}\end{array}\right)|B\in{\rm Sym}_{d}(R)\right\}.$
@@
$E(g,s,\varphi)\coloneqq\sum\limits_{\gamma\in P(\mathbb{Q})\backslash{\mathrm{Sp}}_{2d}(\mathbb{Q})}\Phi(\gamma g,s,\varphi)\quad g\in{\rm Mp}_{2d}(\mathbb{A}),\quad s\in\mathbb{C},$
@@
The famous Siegel–Weil formula identifies the theta lifting of the constant function $\mathbf{1}$ and the special value of the Eisenstein series.
Theorem A.4 (Siegel–Weil formula in the metaplectic case (cf. [MpSW])).
@@
Suppose $M$ is a positive definite even lattice of rank $r$ with $\frac{r}{2}>d+1$. Then
@@
The idea is to identify the component functions of both sides via the Siegel–Weil formula. When $r$ is even, this is [Ma24, Thm. 5.5]. Now suppose $r$ is odd and we sketch the proof as below.
According to Kudla (cf. [Kudla96, p. 37 Proposition 4.3]), for the archimedean component of the Weil representation $\omega_{\psi}$, there is a character $\chi_{V,\infty}^{\psi}$ of $\widetilde{A(\mathbb{R})}$ such that
$(\omega_{\psi,\infty}(\tilde{m})\varphi_{\infty})(\mathbf{x})=\chi_{V,\infty}^{\psi}(\tilde{m})|\det(m(a))|^{\frac{r}{2}}\varphi_{\infty}(\mathbf{x}\cdot a),$
where $\tilde{m}=(m(a),\phi_{m(a)})\in\widetilde{A(\mathbb{R})}$, $a\in{\rm GL}_{d}(\mathbb{R})$ and $\varphi_{\infty}\in\mathcal{S}(V(\mathbb{R})^{d})$.
Let $\varphi_{\infty}(\mathbf{x})=e^{\pi\mathrm{tr}(\mathbf{x},\mathbf{x})}$ be the standard Gaussian function and for $\gamma\in(M^{\vee}/M)^{d}$, denote $\varphi_{\infty}\otimes\varphi_{\gamma}$ by $\varphi_{\infty,\gamma}$. For $\tau=x+iy\in\mathbb{H}_{d}$ and $a\in{\rm GL}_{d}(\mathbb{R})$ satisfying $aa^{\top}=y$, we consider the element
@@
The component function of our previously defined vector-valued genus theta series (resp. vector-valued Siegel–Eisenstein series) is given as follows.
@@
    $\langle\Theta^{(d)}_{\mathbf{Gen}(M)}(\tau),\mathfrak{e}_{\gamma}\rangle\\ =\chi_{V,\infty}^{\psi}(\tilde{m})^{-1}|\det(m(a))|^{-\frac{r}{2}}\int_{\mathrm{O}(V)(\mathbb{Q})\backslash\mathrm{O}(V)(\mathbb{A})}\theta(\widetilde{g}_{\tau},h,\varphi_{\infty,\gamma})\mathrm{d}h,$
    where $\widetilde{g}_{\tau}=n(x)\tilde{m}=(g_{\tau},\phi_{g_{\tau}})$ and the Haar measure $\mathrm{d}h$ is normalized so that
    ${\rm vol}(\mathrm{O}(V)(\mathbb{Q})\backslash\mathrm{O}(V)(\mathbb{A}))=1.$
  • Siegel–Eisenstein series:
@@
For the comparison of theta series, the proof is exactly the same as in the classical case (cf. [LiChao21, Example 2.2.6]). For more details, one can see [Ma24, Thm. 5.5]. For the comparison of Siegel–Eisenstein series, since [Ma24, Thm. 5.5] uses a computation by Kudla which might be difficult to locate a reference in the metaplectic case, we briefly present an alternative proof here.
Consider $q_{\tau}(\mathbf{x})\coloneqq e^{-\pi i\mathrm{tr}((\mathbf{x},\mathbf{x})\tau)}$. One has
@@
for $n(B)=\Big{(}\left(\begin{array}[]{cc}\mathrm{I}_{d}&B\\ 0&\mathrm{I}_{d}\end{array}\right),1\Big{)}$, where $B\in{\rm Sym}_{d}(\mathbb{Z})$ and
@@
for $J_{d}=\Big{(}\left(\begin{array}[]{cc}0&-\mathrm{I}_{d}\\ \mathrm{I}_{d}&0\end{array}\right),\sqrt{\det(\tau)}\Big{)}$. Since ${\rm Mp}_{2d}(\mathbb{Z})$ is generated by $J_{d}$ and $n(B)$’s, combining these equations one gets
@@
for $\tilde{g}=(g,\phi_{g}(\tau))\in{\rm Mp}_{2d}(\mathbb{Z})$. [^2] According to Shimura [Shi85], this is basically how one gets the classical definition of ${\rm Mp}_{2d}(\mathbb{R})$ as in Section 2 via the Weil representation. For details, one can see for example [Hida07, Proposition 2.4, Corollary 2.5]. Hence
@@
and the rest are exactly the same as the last several lines in [Ma24, Thm. 5.5].
@@ Appendix B Coset decomposition and Weil representations
It has been shown in [Ma24, Proposition 4.2] that for any $g^{(2)}\in{\mathrm{Sp}}_{4}(\mathbb{Z})$ there exist some $\alpha\in\mathbb{Z}^{\leq 0},A,B\in{\rm SL}_{2}(\mathbb{Z})$ such that
@@
and the map
$\varphi:{\mathrm{Sp}}_{4}(\mathbb{Z})$ $\to\mathrm{Mat}_{2}(\mathbb{Z}),$
$\varphi(D\mathcal{C}_{\alpha}u(A)d(B))$ =±$=\pm B^{\prime}\begin{pmatrix}\alpha^{2}&0\\ 0&1\end{pmatrix}A,$
@@ 2. $(ii)$
$\phi_{(A\cdot B)^{\prime}}(z)\sqrt{z^{\prime}+(A\cdot B)^{\prime}z}$ $=\phi_{A^{\prime}}(z)\phi_{B}(z^{\prime})\sqrt{A^{\prime}z+Bz^{\prime}}$
$=\phi_{B^{\prime}\cdot A^{\prime}}(z)\sqrt{z^{\prime}+(A\cdot B)^{\prime}z}.$
@@
This is clear because we have
@@
Now it follows from [Ma24, Proposition 4.2] that there is a well-defined injection
@@
Lemma B.3 ([Ma24, Lem. 4.3]).
Let $\alpha\in\mathbb{Z}$, $\tilde{A}\in\widetilde{\mathbf{Y}}_{\alpha^{2}}$ and $\tilde{B}\in{\rm Mp}_{2}(\mathbb{Z})$. Then
@@
Given this decomposition, when $\tilde{g}^{(2)}=\widetilde{\mathcal{C}}_{\alpha}$, the equation B.5 holds by a direct computation (cf. [Ma24, Proposition 4.4]). [^3] A careful reader may note that since our definition of ”$J$” is different from that in [Ma24], the action of ”$J_{2}$” in [Ma24] is actually the action of ”$J_{2}^{-1}$” in our notation. Therefore, the reason why we use $J_{2}^{-1}$ in the above decomposition is to keep the consistency with the computations in [Ma24, Proposition 4.4]. Then from the decomposition (B.1), we have
$\rho_{M}^{(2)}$ $(\tilde{g}^{(2)})^{-1}(\mathfrak{e}_{0}\otimes\mathfrak{e}_{0})$
$=\rho_{M}^{(2)}(\widetilde{\mathcal{C}}_{\alpha}u(\tilde{A})d(\tilde{B}))^{-1}(\mathfrak{e}_{0}\otimes\mathfrak{e}_{0})$
@@ Appendix C Hecke kernels
Denote the element $\left(T=\left(\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right),1\right)\in{\rm Mp}_{2}(\mathbb{Z})$ by $\widetilde{T}$. Then $K_{1}(z,z^{\prime})$ equals
$\quad\sum\limits_{\gamma\in G_{M}}\sum\limits_{\tilde{g}\in{\rm Mp}_{2}(\mathbb{Z})}\frac{\phi_{g}(z)^{-2k}}{(z^{\prime}+g\cdot z)^{k}}\rho_{M}(\tilde{g})^{-1}(\mathfrak{e}_{\gamma})\otimes\mathfrak{e}_{\gamma}$ (C.2)
@@
where $\mathbf{P}_{\gamma,m}(z)=\frac{1}{2}\sum\limits_{\tilde{g}\in\widetilde{\Gamma}_{\infty}^{+}\backslash{\rm Mp}_{2}(\mathbb{Z})}(\mathbf{e}(mz)\mathfrak{e}_{\gamma})\mid_{k,G_{M}}[\tilde{g}]$ is the Poincaré series defined in [Br02, §1.2]. Here, the last identity comes from the Lipschitz Summation Formula in [Lip89] (see also [PPA01, Thm. 1]), which we recall in Lemma C.2 below.
@@
By [Br02, Proposition 1.5], one has $\langle f,\mathbf{P}_{\gamma,m}(z)\rangle_{\rm Pet}=\frac{2\cdot\Gamma(k-1)}{(4\pi m)^{k-1}}c(m,\gamma)$. Combine this with (C.2), we finally get
@@
Acknowledgements
We thank Jan Hendrik Bruinier, Kaiyuan Gu, Xuanlin Huang, Kai-Wen Lan, and Fei Si for helpful conversations and suggestions, and Xinyi Yuan for his interest in this project. We are also grateful to the two anonymous referees for their insightful comments, which helped to improve the introduction of the paper.